OpenStudy (anonymous):

The sum of three consecutive integers is 72. Find the integers. help please?

OpenStudy (anonymous):

1st integer = x 2nd integer = x + 1 3rd integer = x + 2 the sum of 3 consecutive integers is 72 x + x + 1 + x + 2 = 72 (combine like terms and solve for x.